# The Coding Limits of Robust Watermarking for Generative Models

### v1

@@ 3.4 From Tamper Detection to Error Correction
Suppose, for contradiction, that there exists a secret-key multi-message code $\Gamma$ achieving $\mathcal{F}$-error correction above the threshold. Fix any message $\mu^{*}$. Define a messageless code $\Gamma^{\prime}$ by $\Gamma^{\prime}.\mathsf{Enc}(\mathit{sk})=\Gamma.\mathsf{Enc}(\mathit{sk},\mu^{*})$ and $\Gamma^{\prime}.\mathsf{Dec}(\mathit{sk},\gamma)=\mathtt{valid}$ if $\Gamma.\mathsf{Dec}(\mathit{sk},\gamma)=\mu^{*}$, and $\mathtt{invalid}$ otherwise. The error correction property for $\Gamma$ ensures tamper detection for $\Gamma^{\prime}$, contradicting the impossibility for messageless codes. For the multi-message setting, the construction of Christ and Gunn [CG24] can be applied to lift any single-message PRC to a multi-message PRC, preserving all security properties. ∎
@@
No secret-key code (including PRCs [CG24]) over a binary alphabet can achieve error correction at tampering rates exceeding $(1+\delta)/2$, for any constant $\delta>0$, without losing soundness.
Thus, the impossibility is fundamental: it holds for all cryptographically meaningful coding schemes, regardless of whether they detect or correct errors, or how many messages they support.
3.5 A Simple Construction
To demonstrate that the tightness of the impossibility threshold for robust watermarking, we give a simple information-theoretic construction of a messageless secret-key code that attains optimal soundness and tamper detection up to the threshold. This construction is explicit and efficient, with all operations linear in the codeword length.
The main idea is to use the secret key itself as the codeword, and have the decoder distinguish between “valid,” “tampered,” and “invalid” based on Hamming distance from the key. This divides the space of possible codewords into three regions, separated by a threshold $t$.
@@ 3.5 A Simple Construction
Sample a secret key $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\Sigma^{n}$ uniformly at random, where $|\Sigma|=q=n^{c}$ for constant $c>1$.
@@
Let $n\in\omega(\log\lambda)$. The messageless secret-key code $\Gamma$ from 1 satisfies soundness and $\mathcal{F}_{\alpha n}$-tamper detection for $\alpha=(1-\frac{1}{q})(1-\delta)$ for any $\delta\in(0,1)$.
@@
Soundness: Fix any string $\hat{\gamma}\in\Sigma^{n}$ and random key $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\Sigma^{n}$. Let $X$ be the number of positions where $\hat{\gamma}[i]=\mathit{sk}[i]$; $X$ is binomial with mean $n/q$. The decoder outputs $\mathtt{invalid}$ unless $X>n(1+\delta)/q$ (note that the last inequality holds whenever $q\geq 2$). [^1] Indeed, since $X$ is the random variable representing the number of identical symbols, the number of differing symbols is $n-X\leq n-n(1+\delta)/q=n\mathopen{}\mathclose{{\left(1-\frac{1}{q}}}\right)(1+\delta)=t$, which is exactly the condition under which the decoder outputs either $\mathtt{tampered}$ or $\mathtt{valid}$.
By Chernoff bound,
$\mathbb{P}\mathopen{}\mathclose{{\left[X\geq\frac{n}{q}(1+\delta)}}\right]\leq\exp\mathopen{}\mathclose{{\left(-\frac{\delta^{2}n}{2q}}}\right)$
which is negligible for $n\in\omega(\log\lambda)$, $q=n^{c}$ (for some constant $c>1$) and constant $\delta>0$.
@@
Achieving Pseudorandomness.
We can upgrade any deterministic messageless secret-key code to achieve pseudorandomness as follows:
Construction 2 (Pseudorandom Messageless Secret-Key Code with Public Counter).
Let $\Gamma=(\mathsf{KGen},\mathsf{Enc},\mathsf{Dec})$ be any messageless secret-key code with codeword length $n$. Let $\mathsf{F}:\{0,1\}^{\lambda}\times\{0,1\}^{p}\to\{0,1\}^{n\log q}$ be a secure pseudorandom function. Define the upgraded code $\Gamma^{\prime}=(\mathsf{KGen}^{\prime},\mathsf{Enc}^{\prime},\mathsf{Dec}^{\prime})$:
Key Generation $\mathsf{KGen}^{\prime}(\lambda)$:
Sample $(\mathit{sk},\kappa){\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\Gamma.\mathsf{KGen}(\lambda)\times\{0,1\}^{\lambda}$.
Encoding $\mathsf{Enc}^{\prime}((\mathit{sk},\kappa),\pi)$:
Output $\gamma^{\prime}=\Gamma.\mathsf{Enc}(\mathit{sk})\oplus\mathsf{F}(\kappa,\pi)$, where $\pi\in\{0,1\}^{p}$ is a public counter, where $\Gamma.\mathsf{Enc}(\mathit{sk})$ is binary.
Decoding $\mathsf{Dec}^{\prime}((\mathit{sk},\kappa),\gamma^{\prime},\pi)$:
Compute $\gamma=\gamma^{\prime}\oplus\mathsf{F}(\kappa,\pi)$ and output $\Gamma.\mathsf{Dec}(\mathit{sk},\gamma)$.
Requirements on the Counter.
To claim security, the value $\pi$ must satisfy the following three conditions: $(i)$ Publicly known to both encoder and decoder, $(ii)$ never reused (each $\pi$ is used at most once), and $(iii)$ not settable or rewritable by the adversary.
Any deterministic, sound, and tamper-detecting code $\Gamma$ yields a pseudorandom code $\Gamma^{\prime}$ as above, provided a secure PRF and a suitable counter implementation. In practice, such a counter may be realized as a blockchain index, a hardware monotonic counter, or a database sequence number, as long as it is monotonic and trusted. All security properties are preserved, and codewords are now pseudorandom to any party not holding $\kappa$.
4 Connecting Watermarking and Tamper Detection
The fundamental limits we have established for messageless secret-key codes also govern the robustness of watermarking schemes for generative models. To make this connection precise, we show how any watermarking scheme that is sound and robust against tampering can be used to build a secret-key code with exactly the same security guarantees.
Theorem 5.
Assume there exists a watermarking scheme $\Pi$ for a class of generative models $\mathcal{M}=\{\mathsf{Model}:\Sigma^{*}\rightarrow\Sigma^{*}\}$ satisfying soundness and $^\hat{\mathcal{F}}$-robustness (for some family $\hat{\mathcal{F}}$). [^2] For the sake of clarity, we assume the prompt and output space of $\mathsf{Model}$ is defined over the same alphabet. The result generalizes to different alphabets as well. Then, there exists a messageless secret-key code $\Gamma$ with alphabet space $\Sigma$} and codeword space $\Sigma$$^{*}$ satisfying soundness and $^\hat{\mathcal{F}}$-tamper detection.
Proof.
We start by describing the messageless secret-key code $\Gamma$. Let $\overline{x}\in\Sigma^{*}$ be a fixed prompt.
Key generation $\Gamma.\mathsf{KGen}({1^{\lambda}})$:
The key-generation algorithm outputs $\mathit{sk}=(\kappa,\mathsf{Model}_{\kappa})$ where $(\kappa,\mathsf{Model}_{\kappa}){\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\Pi.\mathsf{Watermark}(\mathsf{Model})$.
Encoding $\Gamma.\mathsf{Enc}(\mathit{sk})$:
The encoding algorithm outputs $\gamma=y\in\Sigma^{*}$ where $y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Model}_{\kappa}(\overline{x})$.
Decoding $\Gamma.\mathsf{Dec}(\mathit{sk},\gamma)$:
where $\tilde{\gamma}\leftarrow\hat{f}(\gamma)$ and where the probability is over the choice of the secret key and the randomness used to generate $\gamma{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Model}_{\kappa}(\overline{x})$. This violates robustness of the watermarking scheme. ∎
When the output of the generative model is a binary string (i.e., ${\Sigma}=\{0,1\}$), the connection above becomes especially powerful. By mapping any robust watermarking scheme to a messageless secret-key code, we can directly apply our main impossibility result for codes to watermarking itself.
Specifically, our earlier theorem (2) shows that no messageless secret-key code can be sound and robust to tampering beyond a certain threshold—even when the codewords may have variable length. By combining this with the reduction above, we obtain an immediate impossibility for robust watermarking of binary outputs:
Corollary 3.
Let $\mathcal{M}=\{\mathsf{Model}:\mathcal{X}\rightarrow\{0,1\}^{*}\}$ be any class of generative models with binary outputs. There is no watermarking scheme for $\mathcal{M}$ that achieves both soundness and $\hat{\mathcal{F}}^{\sf ind}_{\alpha}$-robustness for any tampering rate $\alpha\geq(1+\delta)/2$, for any $\delta\in(0,1)$. This holds even if the model’s prompt space $\mathcal{X}$ is arbitrary, and even if the soundness is relaxed to allow a constant error probability.
The impossibility threshold we have identified for tamper-resilient codes is not just a technical detail; it serves as a universal limit for watermarking. Any watermarking scheme for generative models with binary outputs must face this same upper bound, no matter how the scheme is constructed or how input and output lengths are chosen. The recent work of Christ and Gunn [CG24] provides a compelling example: their PRC-based watermarking schemes achieve robustness against independent bit-flip tampering up to the threshold $\alpha<(1-\delta)/2$, and their approach even accommodates certain types of deletions. Our corollary shows that this rate is not only achievable but also optimal. Although we do not claim a perfect correspondence between all watermarking schemes and all PRC constructions, since this depends on the precise model and tampering family, the upper bound on robustness holds for all such schemes.
5 A Concrete Attack
5.1 Setup
Our goal is to test whether the theoretical robustness threshold we proved manifests in practice. For this purpose, we analyze the PRC watermark of Gunn, Zhao, and Song [GZS25], implemented in Stable Diffusion 2.1 Base ($\times512$ resolution, 50 denoising steps). This watermark is representative of the state of the art: it preserves perceptual quality, is provably undetectable under standard assumptions, and resists a wide class of natural manipulations.
[…]
5.2 First Attempts
Before developing our attack, we tested the robustness of [GZS25] while preserving image quality, by chosing a broad set of common image manipulations (including gaussian noise, gaussian blur, pixel-wise color shifts, lossy format conversions).
Image manipulations.
Pixel-wise color shifts.
Lossy format conversions.
(a) Original Image
(b) Color Attack: 23% Error
(c) HSV Attack: 26% Error
(d) JPEG Attack: 32% Error
(e) WebP Attack: 34% Error
Representative examples of these attacks are shown in Figure 1. The images have been visibly altered (sometimes severely) yet the watermark survived. The lesson is clear: edits that preserve image quality produce far too few sign flips to reach the $50\%$ threshold, while edits that come closer to the threshold do so only at the cost of destroying the image. To break the watermark without sacrificing fidelity, a qualitatively different kind of transformation is required.
5.3 Our Successful Attack
(a) Original
(b) Cropped & Resized
(c) Original
(d) Cropped & Resized
(e) Original
(f) Cropped & Resized
(g) Original
(h) Cropped & Resized
5.4 Why It Works
Latent-space effects.
Figure 3 shows this effect in detail. The original latent sign pattern (top left) encodes the watermark (i.e., the PRC codeword) as a pseudorandom arrangement of signs. After crop-and-resize, the recovered latent (bottom left) is visibly scrambled. The difference map before decoding (top right) shows that $48.07\%$ of the signs have flipped. This is essentially the information-theoretic boundary: once errors approach $50\%$, the codeword is indistinguishable from random. This is exactly the core argument to prove the impossibility result (proven by demonstrating a generic attack) of 2. The bottom-right panel shows the result after belief propagation, the error-correcting decoder, has attempted to repair the errors. At this level of corruption, BP cannot recover the original codeword and instead converges to a different pseudorandom codeword. As a result, the measured error rises slightly to $48.96\%$. This increase is not paradoxical but diagnostic: the decoder has lost all correlation with the true codeword. The image itself is unchanged, but the watermark is irretrievable.
(a) Original Image
(b) Crop & Resize (47.2% error)
(c) Crop & Pad (16.7% error)
(d) Downscale & Pad (49.8% error)
(e) Downscale & Upscale (12.1% error)
(a) Crop & Resize
(b) Crop & Pad
(c) Downscale & Pad
(d) Downscale & Upscale
Variants.
(a) Original Watermarked Image
(b) Watermark Removed Output (Cropping 10px)
Figure 6: Demonstration of watermark removal via cropping attack for prompt generated with border
Generalization.
Acknowledgments

### v2

@@ 4 Connecting Watermarking and Tamper-detection Codes
We start by describing the zero-bit tamper-detection code $\Gamma$. Let $\overline{x}\in\Sigma^{*}$ be a fixed prompt.
Key generation $\Gamma.\mathsf{KGen}({1^{\lambda}})$:
The key-generation algorithm outputs $\mathit{sk}=(\kappa,\mathsf{Model}_{\kappa})$ where $(\kappa,\mathsf{Model}_{\kappa}){\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\Pi.\mathsf{Watermark}(\mathsf{Model})$.
Encoding $\Gamma.\mathsf{Enc}(\mathit{sk})$:
The encoding algorithm outputs $\gamma=y\in\Sigma^{*}$ where $y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Model}_{\kappa}(\overline{x})$.
Decoding $\Gamma.\mathsf{Dec}(\mathit{sk},\gamma)$:
where $\tilde{\gamma}\leftarrow\hat{f}(\gamma)$ and where the probability is over the choice of the secret key and the randomness used to generate $\gamma{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Model}_{\kappa}(\overline{x})$. This violates robustness of the watermarking scheme. ∎
When the output of the generative model is a binary string (i.e., ${\Sigma}=\{0,1\}$), the connection above becomes especially powerful. By mapping any robust watermarking scheme to a zero-bit tamper-detection code, we can directly apply our main impossibility result for codes to watermarking itself.
Specifically, our earlier theorem (1) shows that no zero-bit tamper-detection code can be sound and robust to tampering beyond a certain threshold—even when the codewords may have variable length or considering constant-bounded soundness. By combining this with the reduction above, we obtain an immediate impossibility for robust watermarking of binary outputs:
@@
Let $\mathcal{M}=\{\mathsf{Model}:\mathcal{X}\rightarrow\{0,1\}^{*}\}$ be any class of generative models with binary outputs. There is no watermarking scheme for $\mathcal{M}$ that achieves both soundness and $\hat{\mathcal{F}}^{\mathrm{ind}}_{\alpha}$-robustness for any tampering rate $\alpha\geq(1+\delta)/2$, for any $\delta\in(0,1)$. This holds even if the model’s prompt space $\mathcal{X}$ is arbitrary, and even when weakening soundness (via an analogous constant-bounded soundness notion to 7) by bounding it by any constant $c\in(0,1]$ (e.g., 1/21/2). [^3] As described for zero-bit tamper-detection codes, this means that the impossibility result rules out even the construction of a “useless” watermarking scheme whose soundness is achieved only with a high constant probability (such as $c=1/2$).
The impossibility threshold we have identified for tamper-resilient codes is not just a technical detail; it serves as a universal limit for watermarking. Any watermarking scheme for generative models with binary outputs must face this same upper bound, no matter how the scheme is constructed or how input and output lengths are chosen. The recent work of Christ and Gunn [C:ChrGun24] provides a compelling example: their PRC-based watermarking schemes achieve robustness against independent bit-flip tampering up to the threshold $\alpha<(1-\delta)/2$, and their approach even accommodates certain types of deletions. Our corollary shows that this rate is not only achievable but also optimal. Although we do not claim a perfect correspondence between all watermarking schemes and all PRC constructions, since this depends on the precise model and tampering family, the upper bound on robustness holds for all such schemes.
5 A Concrete Attack
(a) Original Image
(b) Color Attack: 23% Error
(c) HSV Attack: 26% Error
(d) JPEG Attack: 32% Error
(e) WebP Attack: 34% Error
5.1 Setup
Our goal is to test whether the theoretical robustness threshold we proved manifests in practice. For this purpose, we analyze the PRC watermark of Gunn, Zhao, and Song [gunn2025undetectable], implemented in Stable Diffusion 2.1 Base ($\times512$ resolution, 50 denoising steps). This watermark is representative of the state of the art: it preserves perceptual quality, is provably undetectable under standard assumptions, and resists a wide class of natural manipulations.
[…]
5.2 First Attempts
Before developing our attack, we tested the robustness of [gunn2025undetectable] while preserving image quality, by chosing a broad set of common image manipulations (including gaussian noise, gaussian blur, pixel-wise color shifts, lossy format conversions).
Image manipulations.
Pixel-wise color shifts.
Lossy format conversions.
5.3 Our Successful Attack
(a) Original
(b) Crop & Resize
(c) Original
(d) Crop & Resize
(e) Original
(f) Crop & Resize
(g) Original
(h) Crop & Resize
(a) Original Image
(b) Crop & Resize (47.2% error)
(c) Crop & Pad (16.7% error)
(d) Downscale & Pad (49.8% error)
(e) Downscale & Upscale (12.1% error)
5.4 Why It Works
Latent-space effects.
Figure 3 shows this effect in detail. The original latent sign pattern (top left) encodes the watermark (i.e., the PRC codeword) as a pseudorandom arrangement of signs. After crop-and-resize, the recovered latent (bottom left) is visibly scrambled. The difference map before decoding (top right) shows that $48.07\%$ of the signs have flipped. This is essentially the information-theoretic boundary: once errors approach $50\%$, the codeword is indistinguishable from random. This is exactly the core argument to prove the impossibility result (proven by demonstrating a generic attack) of 1. The bottom-right panel shows the result after belief propagation, the error-correcting decoder, has attempted to repair the errors. At this level of corruption, BP cannot recover the original codeword and instead converges to a different pseudorandom codeword. As a result, the measured error rises slightly to $48.96\%$. This increase is not paradoxical but diagnostic: the decoder has lost all correlation with the true codeword. The image itself is unchanged, but the watermark is irretrievable.
Variants.
(a) Crop & Resize
(b) Crop & Pad
(c) Downscale & Pad
(d) Downscale & Upscale
Generalization.
(a) Original
(b) Cropped (10px)
Figure 6: Watermark removal on bordered image.
Acknowledgments
Appendix A A Construction of Zero-bit Tamper-detection Code
In this appendix we give an explicit information-theoretic construction of a zero-bit tamper-detection code. Our purpose is not to propose a practical scheme, but to show that the robustness threshold from Section 3.2 is tight: there are codes that achieve soundness and tamper detection for any corruption rate strictly below the critical value. The construction is deliberately simple. The secret key itself serves as the codeword, and the decoder classifies a candidate string according to its Hamming distance from the key. This partitions the space of strings into three regions corresponding to $\mathtt{valid}$, $\mathtt{tampered}$, and $\mathtt{invalid}$, separated by a distance threshold $t$.
@@ Appendix A A Construction of Zero-bit Tamper-detection Code
Sample a secret key $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\Sigma^{n}$ uniformly at random, where $|\Sigma|=q\geq 2$.
@@
Let $n$ and $q=|\Sigma|$ such that $\frac{n}{q}\in\omega(\log\lambda)$. [^5] This can always be enforced by increasing the codeword length $n$ with respect to the cardinality $q$ of $\Sigma$. The zero-bit tamper-detection code $\Gamma$ from 1 satisfies soundness and $\mathcal{F}_{\alpha n}$-tamper detection for $\alpha=(1-\frac{1}{q})(1-\delta)$ for any $\delta\in(0,1)$.
@@
Soundness: Fix any string $\hat{\gamma}\in\Sigma^{n}$ and random key $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\Sigma^{n}$. Let $X$ be the number of positions where $\hat{\gamma}[i]=\mathit{sk}[i]$; $X$ is binomial with mean $n/q$. The decoder outputs $\mathtt{invalid}$ unless
$X>n-t=n-n\Bigl(1-\frac{1}{q}\Bigr)(1-\delta)\geq\frac{n}{q}(1+\delta),$
where the last inequality holds whenever $q\geq 2$.
By the Chernoff bound,
$\mathbb{P}\mathopen{}\mathclose{{\left[X\geq\frac{n}{q}(1+\delta)}}\right]\leq\exp\mathopen{}\mathclose{{\left(-\frac{\delta^{2}n}{3q}}}\right),$
which is negligible whenever $\frac{n}{q}\in\omega(\log\lambda)$ and $\delta>0$ is constant.